OpenStudy (anonymous):

If E partitions CD by a ratio of 1/8, then which of the following could be the location of E?

OpenStudy (anonymous):

-7 -5 -3 -1

OpenStudy (whpalmer4):

We're going to need a bit more information about CD!

OpenStudy (anonymous):

ok i attach the line

OpenStudy (anonymous):

\

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

what is the length of CD?

OpenStudy (anonymous):

16?

OpenStudy (anonymous):

@whpalmer4 is that right ?

OpenStudy (whpalmer4):

Let me look!

OpenStudy (anonymous):

ok

OpenStudy (whpalmer4):

yes, that appears to be correct.

OpenStudy (whpalmer4):

are you sure that 1/8 is the ratio?

OpenStudy (anonymous):

yes it is

OpenStudy (whpalmer4):

because if it is, we won't land exactly on any of those numbers in the answer choices

OpenStudy (anonymous):

did you see the picture i attached?

OpenStudy (whpalmer4):

we have 16 units to divide into 1+8=9 parts, and then we'll take 1/9 of those parts from an endpoint to find our partition.

OpenStudy (anonymous):

ok so how do i do that?

OpenStudy (whpalmer4):

possibly the problem expects you to choose -7 as your answer, but that isn't strictly speaking correct. partition CD by 1/8 means that CE and ED have the ratio 1:8, which means we must chop CD into 9 equal parts, and place E 1 part toward D from C. CD goes from -9 to +7 in your diagram, which is a total of 16 units. 16/(8+1) = 16/9 and neither 1*16/9 nor 8*16/9 is a round number, yet all the answer choices are round numbers.

OpenStudy (whpalmer4):

16/9 = 1.77777777... so if we add -9 + 1.777777... we get -7.22222... which is close to -7, but not the same. But all the other answers are even further removed from the 1/8 point, so that's the best of the bunch, I suppose.

OpenStudy (anonymous):

ok that makes sense maybe were supposed to round

OpenStudy (whpalmer4):

If the ratio was 1/7, then we could do the problem correctly. CD is 16 units long, divided into 1+7=8 parts gives us 2 units per part. Our ratio calls for us to go 1 part = 2 units toward D from C, so -9 + 2 = -7 and that is exactly what one of the answer is. That's what made me ask if you had copied the problem correctly (you had, but I think the problem is incorrect. It happens much more often than it should, given what textbook publishers charge!)

OpenStudy (whpalmer4):

Anyhow, do you at least understand how I worked the problem?

OpenStudy (anonymous):

yes i do thank you so much ill put -7 but ill eamil my teacher

OpenStudy (whpalmer4):

Yes, that would be good. If they insist that the problem is correct as written, let me know and I will supply you with some problems for them that demonstrate the flaw in their understanding :-)